# https://GVGEN.github.io

### v1

@@ Abstract
In recent years, 3D Gaussian splatting has emerged as a powerful technique for 3D reconstruction and generation, known for its fast and high-quality rendering capabilities. Nevertheless, these methods often come with limitations, either lacking the ability to produce diverse samples or requiring prolonged inference times. To address these shortcomings, this paper introduces a novel diffusion-based framework, GVGEN, designed to efficiently generate 3D Gaussian representations from text input. We propose two innovative techniques: (1) Structured Volumetric Representation. We first arrange disorganized 3D Gaussian points as a structured form GaussianVolume. This transformation allows the capture of intricate texture details within a volume composed of a fixed number of Gaussians. To better optimize the representation of these details, we propose a unique pruning and densifying method named the Candidate Pool Strategy, enhancing detail fidelity through selective optimization. (2) Coarse-to-fine Generation Pipeline. To simplify the generation of GaussianVolume and empower the model to generate instances with detailed 3D geometry, we propose a coarse-to-fine pipeline. It initially constructs a basic geometric structure, followed by the prediction of complete Gaussian attributes. Our framework, GVGEN, demonstrates superior performance in qualitative and quantitative assessments compared to existing 3D generation methods. Simultaneously, it maintains a fast generation speed ($\sim$7 seconds), effectively striking a balance between quality and efficiency. *** Equal Contribution. $\dagger$ Corresponding Authors.
@@ 1 Introduction
Previous text-to-3D approaches can be broadly classified into two categories: optimization-based generation [33, 48, 8, 41] and feed-forward generation [31, 32, 40, 19]. Optimization-based methods have become rather popular recently, due to the rapid development of text-to-image diffusion models [12, 36]. These methods usually optimize 3D objects conditioned on texts or images through Score Distillation Sampling (SDS)  [33], distilling rich knowledge from 2D image generation models. Despite yielding impressive results, optimization-based methods face the Janus problem [33], manifesting as multiple faces or over-saturation problems. Additionally, the optimization of a single object can be prohibitively time-consuming, requiring extensive computational effort. Contrarily, feed-forward approaches strive to generate 3D assets directly from text descriptions, thus sidestepping the Janus problem and significantly hastening the generation process. Our work is closely related to feed-forward-based methods. However, feed-forward methods that utilize multi-view generation models often create lower-resolution 3D assets than their multi-view image counterparts. Moreover, models directly generating 3D objects from texts often encounter difficulties with semantics when complex prompts are used.
Different from previous feed-forward-based methods like [19] that follow a text-2D-3D framework, our method proposes to generate 3D assets via directly learning 3D representation. In this study, we introduce an innovative and streamlined coarse-to-fine generation pipeline, GVGEN, for generating 3D Gaussians directly from text descriptions. Leveraging the highly expressive and fast-rendering capabilities of 3D Gaussians, our method achieves not only promising results but also maintains rapid text-to-3D generation and rendering. As shown in Fig. 1, our method consists of two stages: GaussianVolume fitting and text-to-3D generation. In the first stage, we introduce GaussianVolume, a structured volumetric form composed of 3D Gaussians. Achieving this is challenging due to the sparse and unstructured nature of optimizing original 3D Gaussians. To address this, we introduce a novel Candidate Pool Strategy for pruning and densification. This approach allows for fitting high-quality volumetric representation of Gaussians, rather than unordered Gaussian points, making the generation process more conducive for a diffusion-based framework, as is utilized in the following step.
@@
  • We propose GVGEN, an efficient text-to-3D coarse-to-fine generation pipeline that first generates geometry volume and then predicts detailed 3D Gaussian attributes, better controlling diverse geometry and appearances of generated assets. GVGEN achieves a fast generation speed ($\sim$7 seconds) compared with baseline methods, effectively striking a balance between quality and efficiency.
@@ 2.1 Text-to-3D Generation
As the emergence of large-scale 3D datasets [11, 10], Feed-forward models are trained to reduce the generation time. Point-E [31] and Shap-E [17], trained on millions of 3D assets, generate point clouds and neural radiance fields respectively. 3D VADER [32] trains auto-decoders to fit latent volumes and uses them to train the diffusion model. Similarly, VolumeDiffusion [40] trains an efficient volumetric encoder to produce training data for the diffusion model. Despite these advancements, text-to-3D Gaussian generation remains largely unexplored, primarily due to the complexity of organizing disorganized Gaussians. Our work introduces a coarse-to-fine pipeline to feed-forward generate 3D Gaussians from texts, by first generating the object’s coarse geometry and then predicting its explicit attributes, pioneering in the direct generation of 3D Gaussians from text descriptions.
@@ 2.2 Differentiable 3D Representation
Since the proposal of Neural Radiance Field (NeRF) [28], various differentiable neural rendering methods [2, 30, 5, 3] have emerged, which demonstrate remarkable capabilities in scene reconstruction and novel view synthesis. These methods are also widely used in 3D generation tasks. Instant-NGP [30] employs feature volumes for accelerations by querying features only at the corresponding spatial positions. Works [5, 3] decompose features into lower dimensions for faster training and less memory storage. However, they still have some shortcomings compared with point-based rendering methods [45, 4] in the aspects of rendering speed and explicit manipulability.
@@ 3.1 Stage1: GaussianVolume Fitting
The original 3D Gaussian Splatting technique offers fast optimization and real-time rendering. However, its lack of structure makes it challenging for 3D neural networks to process effectively. Directly generating 3D Gaussians through networks involves integrating existing point cloud networks, treating Gaussians as carriers of semantically rich features. However, previous work [43] has pointed out the difficulty of directly generating point clouds with only 3-dimensional color features. Extending such methods, especially for high-dimensional Gaussians, may make the learning process very challenging. We also attempted to extend previous work [27] using point cloud diffusion to generate 3D Gaussians, while the results were disappointing.
@@ 3.1.1 GaussianVolume
In 3D Gaussian Splatting [18], features of 3D Gaussians $G$ include the following attributes: a center position $\mu\in\mathbb{R}^{3}$, covariance matrix $\Sigma$, color information $c\in\mathbb{R}^{3}$ (when SH order=0) and opacity $\alpha\in\mathbb{R}$. To better optimize the covariance matrix, $\Sigma$ is analogous to describing the configuration of an ellipsoid via a scaling matrix $S\in\mathbb{R}^{3}$ and a rotation matrix $R\in\mathbb{R}^{3\times 3}$, satisfying:
@@ 3.1.2 Candidate Pool Strategy
We can not directly apply the original interleaved optimization/density strategy to move 3D Gaussians in our approach, due to the dynamical changes in the number of Gaussians not suitable for a fixed number of Gaussians. Densifying or pruning Gaussians freely is not allowed since Gaussians are bijective with assigned grid points. A naive way to avoid the problem is to only adjust the offsets $\Delta\mu$ relying on gradient back-propagation. Unfortunately, we experimentally found that the movement range of Gaussian centers becomes largely limited without a pruning and densification strategy, which leads to lower rendering quality and inaccurate geometry (see Fig. 7(a)). For this reason, we propose a novel strategy (Algorithm 1) to densify and prune the fixed number of Gaussians. The key point to our strategy is storing pruned points in a candidate pool $P$ for later densification.
We initially align Gaussian centers, $\mu$, with assigned positions, $p$ and set offsets $\Delta\mu=0$ (Line 2 in Algorithm 1). The candidate pool $P$ is initialized as an empty set (Line 3). This pool stores "deactivated" points, which refer to pruned points during optimization – they are not involved in the forward and backward process. We optimize each asset for $T$ iterations. Following original 3D Gaussian Splatting [18], thresholds $\tau_{p},\tau_{d}$ on view-space position gradients are used for pruning and densifying every fixed iterations (IsRefinementIteration(t) in Line 8).
@@ 4.1.1 Baseline Methods
We compare with both feed-forward-based methods and optimization-based methods. In the feed-forward category, we evaluate Shap-E [17] and VolumeDiffusion [40], both of which directly generate 3D assets. For optimization-based methods, we consider DreamGaussian [39] as the baseline, where coarse 3D Gaussians undergo optimization with SDS loss [33] and are subsequently converted into meshes for further refinement. For VolumeDiffusion, we report results generated in the feed-forward process, without post-optimization using SDS via pretrained text-to-image models, to ensure fair comparisons with other methods.
@@ 4.1.2 Dataset
Our training dataset comprises the Objaverse-LVIS dataset [11], which contains $\sim$ 46,000 3D models in 1,156 categories. For text prompts for training, we use captions from Cap3D [26], which leverages BLIP-2 [20] to caption multi-view images of objects and consolidates them into single captions through GPT-4 [1]. For evaluation, we generate 100 assets and render 8 views for each asset. The camera poses for these views are uniformly sampled around the object. Our method excels in both visualization and semantic alignment.
@@ 4.2.2 Text-to-3D Generation
Figure 6: Display of Generation Diversity. (a) displays diverse generated assets according to the same text prompt via GVGEN. (b) shows the comparisons among our image-conditioned GVGEN, OpenLRM [13] and TGS [49]. The single-view reconstruction models suffer from the problem of average patterns, leading to implausible shapes and textures in unseen regions, while GVGEN produces reasonable appearances and geometries.
@@ 4.2.3 Generation Diversity
As depicted in Fig. 6(a), GVGEN can generate diverse assets conditioned on the same prompt. The generative diversity of our method not only differentiates it from reconstruction approaches but also heightens the imaginative capabilities of users. Furthermore, we develop an image-to-3D model conditioned on CLIP image embeddings and compare the results with the recently popular single-image reconstruction models TGS [49] and LRM [14] (See Fig. 6(b)). Since we do not have access to the close-sourced LRM, we use a re-implementation of it, namely OpenLRM [13]. The single-view reconstruction models suffer from the problem of average patterns, leading to implausible shapes and textures in unseen regions, while GVGEN produces reasonable appearances and geometries. Such comparisons accentuate the critical difference between our GVGEN and the reconstruction methods.
Table 1: Quantitative Comparisons against Baseline Methods on CLIP Score and Inference Speed. The metrics demonstrate the superior performance of our method.
@@
Table 2: Quantitative Metrics for Ablation Studies. The left table analyzes GaussianVolume fitting strategies, while the right table compares different losses training the GaussianVolume attributes prediction model. For qualitative comparisons, see Fig. 7.
@@ 4.3.1 GaussianVolume Fitting
We first study the effects of strategies fitting GaussianVolume via visual comparisons. In these ablation experiments, 3D assets are trained with 72 images, with camera poses uniformly distributed at a distance of 2.4 from the world center, and evaluated with 24 images posed uniformly at a distance of 1.6. We set volume resolution $N=32$, i.e. number of Gaussian points $N^{3}=32,768$. The full method, as used in the data preparation stage, performs the best in terms of rendering results, even with only $32,768$ Gaussians. "w/o CPS" refers to optimizing coordinate offsets $\Delta\mu$ only via backpropagation, while "w/o offsets" means fixing Gaussian coordinates $\mu=p$ without optimizing point positions. Fig. 7(a) shows that using the full method produces the best rendering results. Tab. 2 reports the average PSNR, SSIM, and LPIPS scores, supporting the effectiveness of the proposed strategy quantitatively. More results can be found in the supplemental materials.
@@ 4.3.2 Text-to-3D Generation
Due to the computational resources required, it is impractical to evaluate each part of our design with models trained on the full Objaverse-LVIS dataset for our ablation study. Therefore, we train the models on a small subset of Objaverse-LVIS to validate the effectiveness of losses in the stage of predicting GaussianVolume attributes. Specifically, we pick 1,336 assets themed animals from 24 LVIS categories for training. We randomly select $\sim$20of the data from the training data and evaluate quantitative metrics with rendered novel views of each asset. Fig. 7(b) demonstrates that the proposed multi-modal losses produce plausible results with fine-grained details, and quantitative results are listed in Tab. 2. The combination of two types of losses leads to more detailed textures while keeping the entire geometry smooth.
@@ 4.4 Limitations
GVGEN has shown encouraging results in generating 3D objects. However, its performance is constrained when dealing with input texts significantly divergent from the domain of training data, as illustrated in Fig. 12 in the supplementary. Since we need to fit the GaussianVolume per object to prepare training data, it is time-consuming to scale up to millions of object data for better diversity. Additionally, the volume resolution $N$ is set as 32 (For 3D Gaussians, the point number is only $N^{3}$=32,768) to save computational resources, which limits the rendering effects of 3D assets with very complex textures. In the future, we will further explore how to generate higher-quality 3D assets in more challenging scenarios.
@@ 5 Conclusions
In conclusion, this paper explores the feed-forward generation of explicit 3D Gaussians conditioned on texts. We innovatively organize disorganized 3D Gaussian points into a structured volumetric form, termed GaussianVolume, enabling feed-forward generation of 3D Gaussians via a coarse-to-fine generation pipeline. To facilitate the fitting of high-quality GaussianVolumes for training, we propose a novel pruning and densifying strategy, i.e. Candidate Pool Strategy. Our proposed framework, GVGEN, demonstrates remarkable efficiency in generating 3D Gaussians from texts. Experimental results demonstrate the competitive capabilities of GVGEN in both qualitative and quantitative terms. This progress suggests potential extensions of our approach in tackling a broader spectrum of challenges within the field.

### v2

# GVGEN: Text-to-3D Generation with Volumetric Representation

@@ Abstract
In recent years, 3D Gaussian splatting has emerged as a powerful technique for 3D reconstruction and generation, known for its fast and high-quality rendering capabilities. Nevertheless, these methods often come with limitations, either lacking the ability to produce diverse samples or requiring prolonged inference times. To address these shortcomings, this paper introduces a novel diffusion-based framework, GVGEN, designed to efficiently generate 3D Gaussian representations from text input. We propose two innovative techniques: (1) Structured Volumetric Representation. We first arrange disorganized 3D Gaussian points as a structured form GaussianVolume. This transformation allows the capture of intricate texture details within a volume composed of a fixed number of Gaussians. To better optimize the representation of these details, we propose a unique pruning and densifying method named the Candidate Pool Strategy, enhancing detail fidelity through selective optimization. (2) Coarse-to-fine Generation Pipeline. To simplify the generation of GaussianVolume and empower the model to generate instances with detailed 3D geometry, we propose a coarse-to-fine pipeline. It initially constructs a basic geometric structure, followed by the prediction of complete Gaussian attributes. Our framework, GVGEN, demonstrates superior performance in qualitative and quantitative assessments compared to existing 3D generation methods. Simultaneously, it maintains a fast generation speed ($\sim$7 seconds), effectively striking a balance between quality and efficiency. Our project page is https://gvgen.github.io/. * Equal Contribution. $\dagger$ Corresponding Authors.
@@ 1 Introduction
Previous text-to-3D approaches can be broadly classified into two categories: optimization-based generation [35, 50, 9, 43] and feed-forward generation [33, 34, 42, 20, 25]. Optimization-based methods have become rather popular recently, due to the rapid development of text-to-image diffusion models [13, 38]. These methods usually optimize 3D objects conditioned on texts or images through Score Distillation Sampling (SDS)  [35], distilling rich knowledge from 2D image generation models. Despite yielding impressive results, optimization-based methods face the Janus problem [35], manifesting as multiple faces or over-saturation problems. Additionally, the optimization of a single object can be prohibitively time-consuming, requiring extensive computational effort. Contrarily, feed-forward approaches strive to generate 3D assets directly from text descriptions, thus sidestepping the Janus problem and significantly hastening the generation process. Our work is closely related to feed-forward-based methods. However, feed-forward methods that utilize multi-view generation models often create lower-resolution 3D assets than their multi-view image counterparts. Moreover, models directly generating 3D objects from texts often encounter difficulties with semantics when complex prompts are used.
Different from previous feed-forward-based methods like [20] that follow a text-2D-3D framework, our method proposes to generate 3D assets via directly learning 3D representation. In this study, we introduce an innovative and streamlined coarse-to-fine generation pipeline, GVGEN, for generating 3D Gaussians directly from text descriptions. Leveraging the highly expressive and fast-rendering capabilities of 3D Gaussians, our method achieves not only promising results but also maintains rapid text-to-3D generation and rendering. As shown in Fig. 1, our method consists of two stages: GaussianVolume fitting and text-to-3D generation. In the first stage, we introduce GaussianVolume, a structured volumetric form composed of 3D Gaussians. Achieving this is challenging due to the sparse and unstructured nature of optimizing original Gaussians. To address this, we introduce a novel Candidate Pool Strategy for pruning and densification. The approach allows for fitting high-quality volumetric representation of Gaussians, rather than unordered points, making the generation process more conducive for a diffusion-based framework, as is utilized in the following step.
@@
  • We propose GVGEN, an efficient text-to-3D coarse-to-fine generation pipeline that first generates geometry volume and then predicts detailed Gaussian attributes, better controlling diverse geometry and appearances of generated assets. GVGEN achieves a fast speed ($\sim$7 seconds) compared with baseline methods, effectively striking a balance between quality and efficiency.
@@ 2.1 Text-to-3D Generation
As the emergence of large-scale 3D datasets [12, 11], Feed-forward models are trained to reduce the generation time. Point-E [33] and Shap-E [18], trained on millions of 3D assets, generate point clouds and neural radiance fields respectively. 3D VADER [34] trains auto-decoders to fit latent volumes and uses them to train the diffusion model. Similarly, VolumeDiffusion [42] trains an efficient volumetric encoder to produce training data. Despite these advancements, text-to-3D Gaussian generation remains largely unexplored, primarily due to the complexity of organizing disorganized Gaussians. Our work introduces a coarse-to-fine pipeline to feed-forward generate 3D Gaussians from texts, by first generating the object’s coarse geometry and then predicting its explicit attributes, pioneering in the direct generation of 3D Gaussians from text descriptions.
@@ 2.2 Differentiable 3D Representation
Since the proposal of Neural Radiance Field [30], various differentiable neural rendering methods [2, 32, 5, 3] have emerged, demonstrating remarkable capabilities in scene reconstruction, novel view synthesis, and 3D generation tasks. Instant-NGP [32] employs feature volumes for accelerations by querying features only at the corresponding spatial positions. Works [5, 3] decompose features into lower dimensions for faster training and less storage. However, they still have some shortcomings compared with point-based rendering methods [47, 4] in the aspects of rendering speed and explicit manipulability.
@@ 3.1 Stage1: GaussianVolume Fitting
The original 3D Gaussian Splatting offers fast optimization and real-time rendering. However, its lack of structure makes it challenging for 3D neural networks to process effectively. Directly generating 3D Gaussians through networks involves integrating existing point cloud networks, treating Gaussians as carriers of semantically rich features. However, previous work [45] has pointed out the difficulty of directly generating point clouds with only 3-dimensional color features. Extending such methods, especially for high-dimensional Gaussians, may make the learning process very challenging. We also attempted to extend previous work [29] using point cloud diffusion to generate 3D Gaussians, while the results were disappointing.
@@ 3.1.1 GaussianVolume
In 3D Gaussian Splatting [19], features of 3D Gaussians $G$ include: a center position $\mu\in\mathbb{R}^{3}$, covariance matrix $\Sigma$, color information $c\in\mathbb{R}^{3}$ (when SH order=0) and opacity $\alpha\in\mathbb{R}$. To better optimize the covariance matrix, $\Sigma$ is analogous to describing the configuration of an ellipsoid via a scaling matrix $S\in\mathbb{R}^{3}$ and a rotation matrix $R\in\mathbb{R}^{3\times 3}$, satisfying:
@@ 3.1.2 Candidate Pool Strategy
We can not directly apply the original interleaved optimization/density strategy to move 3D Gaussians , due to the dynamical changes in the number of Gaussians not suitable for a fixed number of Gaussians. Densifying or pruning Gaussians freely is not allowed since Gaussians are bijective with assigned grid points. A naive way to avoid the problem is to only adjust the offsets $\Delta\mu$ relying on gradient back-propagation. Unfortunately, we experimentally found that the movement range of Gaussian centers becomes largely limited without an optimization/density strategy, which leads to lower quality and inaccurate geometry (see Fig. 7(a)). For this reason, we propose a novel strategy (Algorithm 1) to densify and prune the fixed number of Gaussians. The key point to our strategy is storing pruned points in a candidate pool $P$ for later densification.
We initially align Gaussian centers, $\mu$, with assigned positions, $p$ and set offsets $\Delta\mu=0$ (Line 2 in Algorithm 1). The candidate pool $P$ is initialized as an empty set (Line 3). This pool stores "deactivated" points, refering to pruned points during optimization – they are not involved in the forward and backward process. We optimize each asset for $T$ iterations. Following original 3DGS [19], thresholds $\tau_{p},\tau_{d}$ on view-space position gradients are used for pruning and densifying every fixed iterations (IsRefinementIteration(t) in Line 8).
@@ 4.1.1 Baseline Methods
We compare with both feed-forward-based methods and optimization-based methods. Across the former methods, we evaluate Shap-E [18] and VolumeDiffusion [42], both of which directly generate 3D assets. For latter ones, we consider DreamGaussian [41] as the baseline, where coarse 3D Gaussians undergo optimization with SDS loss [35] and are converted into meshes for further refinement. For VolumeDiffusion, we report results generated in the feed-forward process, without post-optimization using SDS via pretrained text-to-image models, to ensure fair comparisons with other methods.
@@ 4.1.2 Dataset
Our training dataset comprises the Objaverse-LVIS dataset [12], which contains $\sim$ 46,000 3D models in 1,156 categories. For text prompts for training, we use captions from Cap3D [28], which leverages BLIP-2 [21] to caption multi-view images of objects and consolidates them into single captions through GPT-4 [1]. For evaluation, we generate 100 assets and render 8 views for each asset. The camera poses for these views are uniformly sampled around the object.
@@ 4.2.2 Text-to-3D Generation
Figure 6: Display of Generation Diversity. (a) displays diverse results with the same text prompt via GVGEN. (b) shows the comparisons among our image-conditioned GVGEN, OpenLRM [14] and TGS [52]. The single-view reconstruction models suffer from the problem of average patterns, leading to implausible shapes and textures in unseen regions, while GVGEN produces reasonable appearances and geometries.
@@ 4.2.3 Generation Diversity
As depicted in Fig. 6(a), GVGEN can generate diverse assets conditioned on the same prompt. The generative diversity of our method not only differentiates it from reconstruction approaches but also heightens the imaginative capabilities of users. Furthermore, we develop an image-to-3D model conditioned on CLIP image embeddings and compare the results with the recently popular single-image reconstruction models TGS [52] and LRM [15] (See Fig. 6(b)). We use OpenLRM [14] as an alternative to the close-sourced LRM. The single-view reconstruction models suffer from the problem of average patterns, leading to implausible shapes and textures in unseen regions, while GVGEN produces reasonable appearances and geometries. Such comparisons accentuate the critical difference between GVGEN and reconstruction methods.
Table 1: Quantitative Comparisons against Baseline Methods on CLIP Score and Inference Speed.
@@
Table 2: Quantitative Metrics for Ablation Studies. The left table analyzes GaussianVolume fitting strategies, while the right table compares different losses training the prediction model. For qualitative comparisons, see Fig. 7.
@@ 4.3.1 GaussianVolume Fitting
We first study the effects of strategies fitting GaussianVolume via visual comparisons. In these ablation experiments, 3D assets are trained with 72 images, with camera poses uniformly distributed at a distance of 2.4 from the world center, and evaluated with 24 images posed uniformly at a distance of 1.6. We set volume resolution $N=32$, i.e. number of Gaussian points $N^{3}=32,768$. The full method, as used in the data preparation stage, performs the best in terms of rendering results, even with only $32,768$ Gaussians. "w/o CPS" refers to optimizing coordinate offsets $\Delta\mu$ only via backpropagation, while "w/o offsets" means fixing Gaussian coordinates $\mu=p$ without optimizing point positions. Fig. 7(a) shows that using the full method produces the best rendering results. Tab. 2 reports the average PSNR, SSIM, and LPIPS scores, supporting the effectiveness of the proposed strategy quantitatively.
@@ 4.3.2 Text-to-3D Generation
Due to the computational resources required, it is impractical to evaluate each part of our design with models trained on the full Objaverse-LVIS dataset for our ablation study. Therefore, we train the models on a small subset of Objaverse-LVIS to validate the effectiveness of losses in the stage of predicting GaussianVolume attributes. Specifically, we pick 1,336 assets themed animals from 24 LVIS categories for training.
We randomly select $\sim$20of the data from the training data and evaluate quantitative metrics with rendered novel views. Fig. 7(b) demonstrates that the proposed multi-modal losses produce plausible results with fine-grained details, and quantitative results are listed in Tab. 2. The combination of two types of losses leads to more detailed textures while keeping the geometry smooth.
@@ 4.4 Limitations
GVGEN has shown encouraging results in generating 3D objects. However, its performance is constrained when dealing with input texts significantly divergent from the domain of training data. It is also time-consuming for fitting millions of training data to scale up for better diversity. Additionally, the volume resolution $N$ is set as 32 to save computational resources, which limits the rendering effects of 3D assets with very complex textures. In the future, we will further explore how to generate higher-quality 3D assets in more challenging scenarios.
@@ 5 Conclusions
In conclusion, this paper explores the feed-forward generation of explicit 3D Gaussians conditioned on texts. We innovatively organize disorganized 3D Gaussian points into a structured volumetric form GaussianVolume with a novel pruning and densifying strategy, i.e. Candidate Pool Strategy, enabling feed-forward generation of 3D Gaussians via a coarse-to-fine generation pipeline. Our proposed framework, GVGEN, demonstrates remarkable efficiency in generating 3D Gaussians from texts. Experimental results demonstrate its competitive capabilities. This progress suggests potential extensions of our approach in tackling a broader spectrum of challenges within the field.
5.0.1 Acknowledgments.
This work was done during Xianglong He’s internship at Shanghai Artificial Intelligence Laboratory. The work is supported by the National Key R&D Program of China (No. 2022YFB4701400/4701402, No. 2022ZD 0160102), SSTIC Grant (KJZD20230923115106012, KJZD20230923114916032), and Beijing Key Lab of Networked Multimedia.